Temperature

To convert from °C to K: T(K) = T(°C) + 273.15

To convert from K to °C: T(°C) = T(K) - 273.15

To convert from °C to °F: T(°F) = T(°C) × 9/5 + 32

To convert from °F to °C: T(°C) = T(°F - 32) × 5/9

A sample has a measured temperature of 516.1 °F. What is its temperature in °C?

°C

Check the approach to obtain the correct conversion equation. Use the temperature in °F to find the temperature in °C.

°C
Answer #1

Given that the equation to convert °F to °C is T(°C) = T(°F - 32) × 5/9 = (516.1-32)*5/9 = 268.944 °C